# Cosmological effect of coherent oscillation of ultralight scalar fields in a multicomponent universe

### v1

@@ II The traditional model: Only one rapidly oscillating scalar field and CDM-like features
ϕ˙2=delimited-⟨⟩superscript˙italic-$\langle\dot{\phi}^{2}\rangle=$ $\frac{1}{2}\left[\dot{\phi}_{+}^{2}+\dot{\phi}_{-}^{2}+m^{2}({\phi}_{+}^{2}+{\phi}_{-}^{2})\right]$ (16)
@@
In the limit $m\gg H$, we see that the above ratio tends to zero, specifying an effective equation of state (EoS) of dust. Till now, we have not found the solution for the scale-factor $a(t)$. To get it, we have to actually solve the Friedmann equation.
@@
This is the matter-dominated universe solution. The crucial point about this solution is that only the above matter-dominated solution of the scale-factor satisfies Eq. (12). It is seen from the above calculations when there is only one coherently oscillating scalar field in the universe, oscillating in a time scale which is much shorter than the cosmological time scale, the cosmological system behaves like a dark matter dominated system and only in these case both Eq. (11) and Eq. (12) are simultaneously satisfied. If the scale-factor is different from the form of the scale-factor in pure matter domination, then Eq. (12) will never be satisfied and the whole calculational scheme will become inconsistent.
To end this section, we show that this scheme of calculation, where we have only one scalar field, is consistent and the scale-factor we have obtained. The second Friedmann equation can be written as
@@
We have seen previously that when $m\gg H$ we have $\langle P\rangle\ll\langle\rho\rangle$. In this limit, we can safely neglect the pressure term in the above equation. In this limi,t the above equation is satisfied for the scale-factor expression given in Eq. (25). This completes the discussion on the equivalence of the coherently oscillating scalar field model and a pure matter dominated phase in a spatially flat FLRW spacetime.
  1. 1.
    The separation of mass or time scales, given by $m\gg H$.
  2. 2.
    The energy density $\rho$ is solely coming from a dust solution.
$=0\,,$
$[\ddot{\phi}_{+}-2m\dot{\phi}_{-}+3H(\dot{\phi}_{+}-m\phi_{-})]\sin\alpha$
$\ddot{\phi}_{+}-2m\dot{\phi}_{-}+3H(\dot{\phi}_{+}-m\phi_{-})$ $=$ $0\,,$ (29)
$\ddot{\phi}_{-}+2m\dot{\phi}_{+}+3H(\dot{\phi}_{-}+m\phi_{+})$ $=$ $0\,.$ (30)
If we again segregate the terms in the above equation depending upon the powers of $m$ then we will get back Eq. (11) and Eq. (12), which will only produce a purely dust dominated universe. This point was explained in the first paragraph of this section. Consequently, in the presence of other matter sources, we do not simplify the above solutions and work with Eq. (29) and Eq. (30) as a set of coupled differential equations in ϕ±subscriptitalic-$\phi_{\pm}$. Here we assume that the universe is not purely dust dominated, although $m\gg H$. In the presence of another perfect fluid component and in the presence of spatial curvature, the Friedmann equations are:
$\frac{\ddot{a}}{a}$ $=$ $-\frac{1}{6}\left[\langle\rho+3P\rangle+(\tilde{\rho}+3\tilde{P})\right]\,,$ (32)
where $\tilde{\rho},\tilde{P}$ arise due to another barotropic fluid or due to the presence of another scalar field and $\Lambda$ specifies the cosmological constant. Here $k=0$ or $k=\pm 1$ specifies the uniform curvature of the spatial 3-dimensional hypersurfaces in the FLRW spacetime. In the previous section, we did not have any spatial curvature energy density as it was purely one component system, and there was only one form of energy density that was present. While we generalize the previous ideas to multicomponent fluids, we can include the curvature energy density as it represents just another kind of energy density which can contribute to the dynamics. If the extra matter sector is coming from another barotropic fluid, then we must have an extra equation:
$\dot{\tilde{\rho}}+3H(\tilde{\rho}+\tilde{P})=0\,.$ (33)
$\langle\rho\rangle=\frac{1}{2}\langle\dot{\phi}^{2}+m^{2}\phi^{2}\rangle=$ $\frac{1}{4}(\dot{\phi}_{+}^{2}+\dot{\phi}_{-}^{2})+\frac{m^{2}}{2}({\phi}_{+}^{2}+{\phi}_{-}^{2})$ (34)
$+\frac{m}{2}(\dot{\phi}_{-}\phi_{+}-\dot{\phi}_{+}\phi_{-})\,.$
$+\frac{m}{2}(\dot{\phi}_{-}\phi_{+}-\dot{\phi}_{+}\phi_{-})\,,$
where we have used Eq. (16) and Eq. (17). We can assume the other fluid has an equation of state (EoS) as:
$\tilde{P}=\tilde{\omega}\tilde{\rho}\,.$ (36)
We have four unknowns: $\phi_{+}(t),\phi_{-}(t),a(t),\tilde{\rho}(t)$ and we can use the four equations as Eq. (29), Eq. (30), Eq. (32) and Eq. (33) to solve for the unknowns. The initial conditions must satisfy Eq. (31). One can always calculate the EoS in the scalar field sector as:
$\omega=\frac{\langle P\rangle}{\langle\rho\rangle}$
The behavior of the effective EoS tell us the nature of the effective matter component of the universe.
As was stated before, we will sometimes assume $\tilde{\rho},\tilde{P}$ arising from another scalar field sector. The other scalar field will be assumed to be representing the dynamic dark energy sector. In those cases, one will work purely with the dynamic dark energy sector and set $\Lambda=0$ in Eq. (31). This scalar field can represent either quintessence ($\epsilon=+1$) or phantom ($\epsilon=-1$) like fields, in the presence of a potential $V(\psi)=V_{0}e^{-\lambda\psi}$, where the scalar field equation is:
$\epsilon\ddot{\psi}+3H\epsilon\dot{\psi}+\frac{dV}{d\psi}=0\,.$ (39)
Here $V_{0}$ and $\lambda$ are parameters which specify the dynamic dark energy potential. This form of potential for dynamic dark energy is widely used in the literature Ratra:1987rm ; Ferreira:1997hj ; Sahni:2002kh ; Copeland:2006wr ; Tsujikawa:2013fta . The corresponding energy density and pressure are:
$\tilde{\rho}$ $=$ $\rho_{\psi}=\frac{\epsilon\dot{\psi}^{2}}{2}+V_{0}e^{-\lambda\psi}\,,$ (40)
$\tilde{P}$ $=$ $P_{\psi}=\frac{\epsilon\dot{\psi}^{2}}{2}-V_{0}e^{-\lambda\psi}\,.$ (41)
When we are mainly interested in observing the effect of dynamical dark energy in the background expanding universe, which contains a CDM component and is assumed to be spatially flat, we will set $k=0$ in Eq. (31). On the other hand, if we want to see how the dark sector behaves in a gravitational collapse of a slightly overdense spherical patch, we will assume $k=1$ in Eq. (31). Next, we discuss how one can include non-minimal coupling between the CDM sector and radiation.
III.1 Inclusion of Non-minimal coupling between the CDM sector and radiation
In Ref. PhysRevD.28.1243 Turner extended the theory of coherently oscillating scalar fields to the case where this scalar field sector has some non-minimal coupling with radiation. Due to the presence of this non-minimal coupling, the scalar field undergoing damped oscillations in the harmonic potential may dump energy into the radiation sector. The oscillations get damped because of the friction produced by the expansion of the universe. In Ref. PhysRevD.28.1243 the effect of the non-minimal coupling was dealt in a heuristic manner, as there was no attempt to observe the backreaction effect of the produced radiation field on the oscillation phenomenology. In our method, we can exactly take into account the effect of non-minimal coupling between the CDM sector and the radiation sector.
In the present case, we will have $\tilde{\rho}=\rho_{(R)}$ and $\tilde{P}=P_{(R)}$ where $\rho_{(R)},P_{(R)}$ are the radiation energy density and radiation pressure. Instead of starting with Eq. (3), we now start with the modified scalar field equation as:
$\ddot{\phi}+(3H+\Gamma)\dot{\phi}+m^{2}\phi=0\,.$ (42)
where $\Gamma$ gives the decay rate of the scalar field. This decaying dark matter scheme is appealing because, in general, axion-like particles have a small electromagnetic coupling Raffelt:1987im and hence axions can decay to photons. Suppose the scalar field decays into radiation, then, in the presence of this non-minimal coupling, Eq. (33) gets modified to:
$\dot{\rho}_{(R)}+4H\rho_{(R)}=\Gamma\dot{\phi}^{2}\,,$ (43)
$\dot{\rho}_{(R)}+4H\rho_{(R)}=\Gamma\langle\dot{\phi}^{2}\rangle$
$+\Gamma m(\dot{\phi}_{-}\phi_{+}-\dot{\phi}_{+}\phi_{-})\,,$ (44)
In general, $\Gamma$ can be a constant or a function of $\phi$. To show the feasibility of our method, in this paper, we have only worked with a constant $\Gamma$. If there is no radiation considered in the system, then the system starts from a point where $\rho_{(R)}=0$ and the radiation density builds up. In the case of constant, $\Gamma$ we can have the weak dissipative regime where $\Gamma<3H$ or the strong dissipative regime where $\Gamma>3H$. Because $\Gamma$ is related to the rate at which the dark matter sector interacts with the radiation sector, we work in the weak dissipative regime. For the equivalence between the coherently oscillating scalar field sector and the CDM sector, one must be careful about the initial point and ensure $m\gg 3H+\Gamma$.
III.2 Intricacy of the initial conditions
@@
In this paper, we are considering a manifold $(\mathcal{M},g_{\alpha\beta})$ that is globally hyperbolic, i.e., there exists a global Cauchy surface ($\Sigma$) whose domain of dependence $D(\Sigma)\equiv\mathcal{M}$. From the Cauchy data given on a Cauchy surface $(\Sigma)$, we can construct the entire manifold. If we consider a single homogeneous scalar field that seeds the spacetime and which is oscillating much faster compared to the cosmological time scale, then the Cauchy data would be: $(\Sigma,\leavevmode\nobreak\ h_{ab},\leavevmode\nobreak\ \mathcal{K}_{ab},\leavevmode\nobreak\ \phi(0),\leavevmode\nobreak\ \dot{\phi}(0))$, where $h_{ab},\mathcal{K}_{ab}$ are the induced metric and the second fundamental form on $\Sigma$, respectively. However, one has to choose the data in such a way that it must satisfy the following constraint equations on $\Sigma$:
@@
~$\tilde{\nabla}_{a}(\mathcal{K}^{ab}-h^{ab}\leavevmode\nobreak\ tr(\mathcal{K}))=\langle j^{b}\rangle,$ (48)
@@
where ${}^{(3)}R$ is Ricci scalar of co-dimension one spacelike $\Sigma$, ~$\tilde{\nabla}_{a}$ is the covariant derivative on that surface, $\mathcal{K}^{2}=\mathcal{K}_{ab}\mathcal{K}^{ab}$, $\langle\rho\rangle=-T^{0}_{0}=-G^{0}_{0}$, and the induced momentum density $\langle j_{a}\rangle=T_{\alpha\beta}\leavevmode\nobreak\ n^{\alpha}\leavevmode\nobreak\ e^{\beta}_{a}$ where $n^{\alpha}$ and $e^{\beta}_{a}$ are the timelike normal and spacelike tangent to the surface $\Sigma$, respectively. Here, the angular brackets specify the time average over the time period of scalar field oscillations. It should be noted that in this paper, the roman indices $\{a,b,c,...\}$ are used to represent the spatial indices while the Greek indices $\{\alpha,\beta,\gamma,...\}$ denote the spacetime indices. Now, in our case, the induced metric on the Cauchy surface is:
$h_{ab}=\begin{pmatrix}\frac{a^{2}(0)}{1-kr^{2}}&0&0\\ 0&r^{2}a^{2}(0)&0\\ 0&0&r^{2}\sin^{2}\theta\leavevmode\nobreak\ a^{2}(0)\end{pmatrix},$ (49)
@@
where the $t=0$ characterizes the spacelike property of the Cauchy surface $\Sigma$. Now, we can compute ($\mathcal{K}_{ab}$) and ${}^{(3)}R$ on $\Sigma$. Using these and also using the first constraint Eq. (47), we get the first Friedmann Eq. (31). It can be verified that the second constraint, Eq.(48), is trivially satisfied for the aforementioned induced metric on $\Sigma$. Therefore, here, the Cauchy problem consists of solving the evolution equation of the metric, i.e., the second Friedmann equation (Eq. (32)) and the evolution equation for $\phi(t)$ (Eq. (3)) or equivalently the evolution equations for $\phi_{+}(t)$ and $\phi_{-}(t)$ (Eq. (29) and Eq. (30) respectively) with the Cauchy data $(\Sigma,\leavevmode\nobreak\ h_{ab},\leavevmode\nobreak\ \mathcal{K}_{ab},\leavevmode\nobreak\ \phi(0),\leavevmode\nobreak\ \dot{\phi}(0))$ or equivalently $(\Sigma,\leavevmode\nobreak\ h_{ab},\leavevmode\nobreak\ \mathcal{K}_{ab},\leavevmode\nobreak\ \phi_{\pm}(0),\leavevmode\nobreak\ \dot{\phi}_{\pm}(0))$ that satisfy the two constraint equations mentioned above. It can be shown that the propagation of the constraint equations along time $(t)$ is evident with this construction.
@@
Now, apart from the constraint equation Eq. (47), there is no intrinsic constraint on scalar fields ($\phi_{\pm}(0)$) and their derivatives $(\dot{\phi}_{\pm}(0))$ on $\Sigma$. Therefore, we are free to choose five initial values out of six $(a(0),\leavevmode\nobreak\ H(0),\leavevmode\nobreak\ \phi_{\pm}(0),\leavevmode\nobreak\ \dot{\phi}_{\pm}(0))$, where $a|_{\Sigma},\leavevmode\nobreak\ H|_{\Sigma}$ are related to $h_{ab},\leavevmode\nobreak\ \mathcal{K}_{ab}$, respectively. For a specific model of scalar fields or space-time dynamics, additional constraint equations may arise that must be satisfied on the Cauchy surface. Next, we discuss the requirement of imposing additional constraints on the scalar fields and their derivatives on $\Sigma$ to ensure that the scalar field closely mimics the behavior of cold dark matter (CDM).
The discussions in the last section established the equivalence between the oscillating scalar field model and the CDM sector, and one may for all practical purpose forget about the scalar field sector and assume that the universe consists of a barotropic fluid whose energy density falls as $a(t)^{-3}$ and whose pressure is zero. On the other hand, in a multicomponent universe, one has to test whether such an equivalence still persists, as in this case, it is difficult to give an analytic proof of the equivalence between the scalar field sector and the CDM sector. Even if one assumes that the previous equivalence holds, then also in the presence of non-minimal coupling of the scalar field $\phi$ with another field or fluid, one has to solve for the field $\phi$. We will first see that even in the simplest model of the equivalence, as discussed in the previous section, the initial constraints on $(a(0),\leavevmode\nobreak\ H(0),\leavevmode\nobreak\ \phi_{\pm}(0),\leavevmode\nobreak\ \dot{\phi}_{\pm}(0))$ which lead to the equivalence are nontrivial. To show the intricacies involved, for the time being, we assume that $\tilde{\rho}=\tilde{P}=0$, $k=0$, $\Lambda=0$, and we have a universe purely dominated by the scalar field energy density. We would like to solve the problem discussed in the last section using the equations derived in this section. In such a case, the Friedmann equations are given by Eq. (14) and Eq. (27), where $\langle\rho\rangle$ and $\langle P\rangle$ are given in Eq. (34) and Eq. (35). The initial data $(a(0),\leavevmode\nobreak\ H(0),\leavevmode\nobreak\ \phi_{\pm}(0),\leavevmode\nobreak\ \dot{\phi}_{\pm}(0))$ that satisfy the Friedmann constraint in Eq. (31) is not enough to reproduce the exact equivalence in the limit $m\gg H$ we obtained in the last section. We observe that the initial values of $\phi_{\pm}(0)$ must satisfy the condition in Eq. (19) on $\Sigma$. If this additional constraint is not satisfied with our initial choice of the functions, then we will never get the previous exact results, as those results were obtained when $\phi_{\pm}(0)$ and $\dot{\phi}_{\pm}(0)$ satisfied the aforementioned condition. If we take into account the condition given in Eq. (13) and take the time derivative of $\phi_{\pm}(t)$ (arising from it) then we see that the initial conditions should be such that
If the initial data $(a(0),\leavevmode\nobreak\ H(0),\leavevmode\nobreak\ \phi_{\pm}(0),\leavevmode\nobreak\ \dot{\phi}_{\pm}(0))$ are chosen to satisfy the Friedmann constraint along with the additional constraints mentioned above, then it can be shown that the scalar field component closely mimics the behavior of cold dark matter (CDM), even in the most general case where multiple matter sources are present in the universe and the scalar field is non-minimally coupled to other components.
$\mathcal{E_{+}}(t)$ $=$ $\dot{\phi}_{+}(t)+\frac{3}{2}\leavevmode\nobreak\ H(t)\phi_{+}(t),$ (51)
$\mathcal{E_{-}}(t)$ $=$ $\dot{\phi}_{-}(t)+\frac{3}{2}\leavevmode\nobreak\ H(t)\phi_{-}(t).$ (52)
For these constraints, one can check that they remain zero, i.e., ±$\mathcal{E_{\pm}}(t)=0$, for all times when we impose the initial constraint as given in Eq. (50) and the scale-factor solution is $a(t)\propto(1+At)^{2/3}$ For this initial constraint, one can have a particular solution of the scale-factor, as given above, for which the terms proportional to $m$ in Eq. (29) and Eq. (30) segregate out, and individually (those terms separately) become zero (as it happened in the previous section). These facts suggest that when we have ±$\mathcal{E_{\pm}}(0)=0$ initially on $\Sigma$, we can have the dust-like solution and the constraints propagate in time. If $a(t)$ has any other time dependence, one may not have ±$\mathcal{E_{\pm}}(t)=0$ on $\Sigma$ and in general ±$\mathcal{E_{\pm}}(t)$ will evolve with time [^3] More analytical work is needed to fully reveal the dynamics of the constraint in general. We hope to work out the dynamics of the constraint propagation in a future publication.. Therefore, if ±$\mathcal{E}_{\pm}|_{t=0}=0$, then ±$\mathcal{E}_{\pm}(t)=0$ for all time $t$, i.e., the additional constraint Eq. (50) on the Cauchy surface $\Sigma$ would propagate along time when we have a pure dust dominated universe. If other non-dust-like matter is present along with dust, then the scale-factor $a(t)$ would not have the similar form, and therefore the additional constraint Eq. (50) at the Cauchy surface on the scalar fields would not propagate exactly.
$\left(\tau\frac{d}{dt}\dot{\phi}_{-}+3H\tau\frac{d}{dt}\phi_{-}\right)+2\left(\dot{\phi}_{+}+\frac{3}{2}H\phi_{+}\right)$ $=$ $0$ (54)
In the above equations, $\tau d/dt$ of any function ϕ˙±subscript˙italic-$\dot{\phi}_{\pm}$ or ϕ±subscriptitalic-$\phi_{\pm}$ simply specifies the change of these functions in the time interval $\tau$. We have assumed that these functions change negligibly in the time interval $\tau$, and consequently, the terms in the brackets on the left side of the (left hand side of the) above equations (which contain $\tau$) are negligible, and consequently, the other terms (in the above equations) must also be negligible. In Ref. PhysRevD.44.352 , the above equations become solvable because all the terms in the brackets, appearing in the above equations, individually vanish for all times $a(t)\propto(1+At)^{2/3}$. In the present case, all the bracketed terms in the above equations remain nonzero; they have negligible values because $m\gg H$. The initial conditions play a special role because we have exactly set $\left(\dot{\phi}_{\pm}(0)+\frac{3}{2}H(0)\leavevmode\nobreak\ \phi_{\pm}(0)\right)=0$ at the Cauchy surface, hence initially all the bracketed terms in the above equations are zero. Because $a(t)\neq a_{i}(1+At)^{2/3}$ in our case, the bracketed terms will not remain exactly zero for later times but will pick up negligible values as we have maintained $m\gg H$ for all times. The constraint in the initial condition, as given in Eq. (50), makes the above equations consistent. If we have not assumed $\left(\dot{\phi}_{\pm}(0)+\frac{3}{2}H(0)\leavevmode\nobreak\ \phi_{\pm}(0)\right)\sim 0$ initially, then the initial conditions should have contradicted Eq. (29) and Eq. (30) when $m\gg H$. We can say that when $a(t)\propto(1+At)^{\frac{2}{3}}$ the constraint $\left(\dot{\phi}_{\pm}(0)+\frac{3}{2}H(0)\leavevmode\nobreak\ \phi_{\pm}(0)\right)=0$ exactly propagates along time, whereas in a multicomponent universe, the constraint propagates approximately. This approximation works very well when $m\gg H$.
From the above discussion we can now conclude that if we allow $k\neq 0$ and more matter components and the cosmological constant in the universe where the Friedmann equations are given by Eq. (31) and Eq. (32) and the quantities $\phi_{\pm}(0),\dot{\phi}_{\pm}(0),a(0),\dot{a}(0)$ satisfy the conditions given in Eq. (50) then the equivalence between the oscillating scalar field sector and the CDM sector remains valid as long as $m\gg H$. Even if the new matter components have a non-minimal coupling with the $\phi$, the initial conditions in Eq. (50) and the condition $m\gg H$ play an active role in preserving the aforementioned equivalence.
Before validating our claims, we will specify some important points about the geometrized units used in our work and explain how they influence the choice of initial conditions. In this unit system, quantities such as length, mass, and time share the same dimensionality, expressed in inverse centimeters ($\text{cm}^{-1}$), while the scalar field $\phi$ remains dimensionless. For broader relevance and comparison, we refer to natural units, where energy is commonly expressed in electron volts (eV). In natural units, we consider the mass range of the oscillating scalar field $\phi$ to be $m\sim 10^{-22}\leavevmode\nobreak\ \text{eV}$, which is considered an ultralight scalar field, while the present-day Hubble constant is $H\sim 10^{-33}\leavevmode\nobreak\ \text{eV}$. In geometrized units, we adopt a scalar field mass of $m=1.93\times 10^{-17}\leavevmode\nobreak\ \text{cm}^{-1}$. Using the conversion factor $1\leavevmode\nobreak\ \text{cm}^{-1}=0.19733\times 10^{-4}\leavevmode\nobreak\ \text{eV}$ our chosen $m$ value corresponds to $m=3.8\times 10^{-22}\leavevmode\nobreak\ \text{eV}$ in natural units, consistent with ultralight scalar field dark matter models.
We further adopt the cosmological parameters from the final results of the Planck satellite full mission, Planck:2018vyg which provide the following present-day values of the density parameters of dark energy ($\Lambda$), matter (M) (composed of baryonic + cold dark matter), cold dark matter (CDM), baryonic matter (B) and radiation (R) as: $\Omega_{(\Lambda)_{0}}=0.68$, $\Omega_{(\text{M})_{0}}=0.315,\quad\Omega_{(\text{CDM})_{0}}h^{2}=0.120,\quad\Omega_{(\text{B})_{0}}h^{2}=0.0224,\quad\Omega_{(\text{R})_{0}}h^{2}=4.15\times 10^{-5}$ with the dimensionless Hubble constant given by $h=0.67$. The redshift z is related to the scale-factor $a$ by the standard relation: $\text{z}=\frac{a_{0}}{a}-1$, with current scale-factor $a_{0}=1$. The scale-factor at radiation-matter equality is given by, $a_{\text{eq}}^{\text{(RM)}}=\frac{\Omega_{(\text{R})_{0}}}{\Omega_{(\text{M})_{0}}}=2.8\times 10^{-4}$ corresponding to a redshift of $\text{z}_{\text{eq}}^{\text{(RM)}}=3571$. Similarly, the scale-factor at matter-dark energy equality is, $a_{\text{eq}}^{\text{(M}\Lambda)}=\frac{\Omega_{(\text{M})_{0}}}{\Omega_{(\Lambda)_{0}}}=0.77$ corresponds to a redshift of $\text{z}_{\text{eq}}^{\text{(M}\Lambda)}=0.3$. The subscript ”00” denotes the values at the current epoch of the universe.
The dark energy component in this study is modeled using two different approaches. The first approach employs the cosmological constant $\Lambda$, with a fixed value of $\Lambda=1.08\times 10^{-56}\leavevmode\nobreak\ \text{cm}^{-2}$, derived from the observed dark energy density parameter, and in geometrized unit the vacuum energy density $\rho_{(\Lambda)}=\Lambda$. The second approach introduces a dynamical scalar field $\psi$, which can be a quintessence or a phantom field. In both cases, the scalar field is governed by an exponential potential of the form $V(\psi)=V_{0}e^{-\lambda\psi}$, where the parameter $\lambda=1$, is fixed throughout the work for both field types. The maximum potential value $V_{0}$ is chosen such that the density parameter of the scalar field matches that of the cosmological constant at the present epoch. We set $V_{0}=1.38\times 10^{-56}\leavevmode\nobreak\ \text{cm}^{-2}$ for the quintessence field and $V_{0}=0.91\times 10^{-56}\leavevmode\nobreak\ \text{cm}^{-2}$ for the phantom field.
To handle extremely small numerical values efficiently, we define a reference unit $u=10^{-28}\leavevmode\nobreak\ \text{cm}^{-1}$, which corresponds to $10^{-33}\leavevmode\nobreak\ \text{eV}$ in natural units. All parameters are expressed in terms of $u$ throughout this paper, thus the value of $m=1.93\times 10^{11}\leavevmode\nobreak\ u$, $V_{0}=1.38\leavevmode\nobreak\ u^{2},0.91\leavevmode\nobreak\ u^{2}$ respectively for quintessence and phantom field and $\rho_{(\Lambda)}=\Lambda=1.08\leavevmode\nobreak\ u^{2}$. To solve the scalar field equation Eq. (39), we require two initial conditions. These are chosen as, $\dot{\psi}(0)=0.0001\leavevmode\nobreak\ u$ and $\psi(0)=0.0001$ and are held fixed throughout the analysis. The initial value of the scale factor $a$ depends on the cosmological epoch under consideration. It becomes dimensionless when a flat universe $k=0$ is assumed, and in cases where the curvature constant $k=1$, the scale factor carries the dimension of length. The specific values used for $a$ will be clarified in the relevant sections.
The energy density $\rho(t)$ for any component is related to the density parameter using $\Omega(t)=\rho(t)/\rho_{c}(t)$, where the critical density in geometrized units is given by $\rho_{c}(t)=3H(t)^{2}$. In the case of an oscillating scalar field, as we are using time-averaged energy density, the corresponding density parameter is defined as $\Omega_{(\phi)}(t)=\langle\rho(t)\rangle/\rho_{c}(t)$. At the present epoch, the critical density takes the value $\rho_{c_{0}}=1.60\leavevmode\nobreak\ u^{2}$. Using this, the present-day dark matter density parameter is expressed as, $\Omega_{(\text{CDM})_{0}}=\frac{\rho_{(\text{CDM})_{0}}}{a^{3}_{0}\rho_{c_{0}}}$, and similarly, for baryonic matter and radiation, we have, $\Omega_{(\text{B})_{0}}=\frac{\rho_{(\text{B})_{0}}}{a^{3}_{0}\rho_{c_{0}}},\quad\Omega_{(\text{R})_{0}}=\frac{\rho_{(\text{R})_{0}}}{a^{4}_{0}\rho_{c_{0}}}$. Using these relations, and $a_{0}=1$ we obtain $\rho_{(\text{CDM})_{0}}=0.42\leavevmode\nobreak\ u^{2},\quad\rho_{(\text{B})_{0}}=0.079\leavevmode\nobreak\ u^{2},\quad\rho_{(\text{R})_{0}}=1.4\times 10^{-4}\leavevmode\nobreak\ u^{2}.$ All these choices, as specified in this section, along with the initial conditions and parameter values, are held fixed throughout our calculations unless stated otherwise.
In this section, we will first verify the predictions made in the previous sections at different levels of complexity. This verification serves as proof of our proposal of the general equivalence of the coherently oscillating ultralight scalar field model and the CDM sector in the general case where there are multiple matter components and the system satisfies the condition $m\gg H$. Initially, we will show that the equivalence works perfectly when we have three more components in the universe, these components being radiation, baryonic matter, and dark energy. Next, we will increase the level of complexity and include non-minimal coupling between the CDM sector and radiation in the presence of dark energy and baryonic matter. The latter case distinctly shows the applicability of the solution presented in this paper in more complicated cosmological settings.

### v2

@@ Appendix A The details of the averaging process
$\langle\dot{\phi}^{2}\rangle=$ $\frac{1}{2}\left[\dot{\phi}_{+}^{2}+\dot{\phi}_{-}^{2}+m^{2}({\phi}_{+}^{2}+{\phi}_{-}^{2})\right]$ (51)
@@
In the limit $m\gg H$, we see that the above ratio tends to zero, specifying an effective equation of state (EoS) of dust.
@@
This is the matter-dominated universe solution. The second Friedmann equation can be written as
@@
We have seen previously that when $m\gg H$ we have $\langle P\rangle\ll\langle\rho\rangle$. In this limit, we can safely neglect the pressure term in the above equation. In this limit the above equation is satisfied for the scale-factor expression given in Eq. (19).
Appendix B Formalism and some more details about the new solutions
@@
If we consider a single homogeneous scalar field that seeds the spacetime and which is oscillating much faster compared to the cosmological time scale, then the Cauchy data would be: $(\Sigma,\penalty 10000\ h_{ab},\penalty 10000\ \mathcal{K}_{ab},\penalty 10000\ \phi(0),\penalty 10000\ \dot{\phi}(0))$, where $h_{ab},\mathcal{K}_{ab}$ are the induced metric and the second fundamental form on $\Sigma$, respectively. However, one has to choose the data in such a way that it must satisfy the following constraint equations on $\Sigma$:
@@
$\tilde{\nabla}_{a}(\mathcal{K}^{ab}-h^{ab}\penalty 10000\ tr(\mathcal{K}))=\langle j^{b}\rangle,$ (64)
@@
where $^{(3)}R$ is Ricci scalar of co-dimension one spacelike $\Sigma$, $\tilde{\nabla}_{a}$ is the covariant derivative on that surface, $\mathcal{K}^{2}=\mathcal{K}_{ab}\mathcal{K}^{ab}$, $\langle\rho\rangle=-T^{0}_{0}=-G^{0}_{0}$, and the induced momentum density $\langle j_{a}\rangle=T_{\alpha\beta}\penalty 10000\ n^{\alpha}\penalty 10000\ e^{\beta}_{a}$ where $n^{\alpha}$ and $e^{\beta}_{a}$ are the timelike normal and spacelike tangent to the surface $\Sigma$, respectively. Here, the angular brackets specify the time average over the time period of scalar field oscillations. It should be noted that in this paper, the roman indices $\{a,b,c,$...\} are used to represent the spatial indices while the Greek indices $\{\alpha,\beta,\gamma,$...\} denote the spacetime indices. Now, in our case, the induced metric on the Cauchy surface is:
$h_{ab}=\begin{pmatrix}\frac{a^{2}(0)}{1-kr^{2}}&0&0\\ 0&r^{2}a^{2}(0)&0\\ 0&0&r^{2}\sin^{2}\theta\penalty 10000\ a^{2}(0)\end{pmatrix}$, (65)
@@
where the $t=0$ characterizes the spacelike property of the Cauchy surface $\Sigma$. Now, we can compute ($\mathcal{K}_{ab}$) and $^{(3)}R$ on $\Sigma$. Using these and also using the first constraint Eq. (63), we get the first Friedmann Eq. (24). It can be verified that the second constraint, Eq.(64), is trivially satisfied for the aforementioned induced metric on $\Sigma$. Therefore, here, the Cauchy problem consists of solving the evolution equation of the metric, i.e., the second Friedmann equation (Eq. (25)) and the evolution equation for $\phi(t)$ (Eq. (3)) or equivalently the evolution equations for $\phi_{+}(t)$ and $\phi_{-}(t)$ (Eq. (22) and Eq. (23) respectively) with the Cauchy data $(\Sigma,\penalty 10000\ h_{ab},\penalty 10000\ \mathcal{K}_{ab},\penalty 10000\ \phi(0),\penalty 10000\ \dot{\phi}(0))$ or equivalently $(\Sigma,\penalty 10000\ h_{ab},\penalty 10000\ \mathcal{K}_{ab},\penalty 10000\ \phi_{\pm}(0),\penalty 10000\ \dot{\phi}_{\pm}(0))$ that satisfy the two constraint equations mentioned above. It can be shown that the propagation of the constraint equations along time $(t)$ is evident with this construction.
@@
Now, apart from the constraint equation Eq. (63), there is no intrinsic constraint on scalar fields ($\phi_{\pm}(0)$) and their derivatives $(\dot{\phi}_{\pm}(0))$ on $\Sigma$. Therefore, we are free to choose five initial values out of six $(a(0),\penalty 10000\ H(0),\penalty 10000\ \phi_{\pm}(0),\penalty 10000\ \dot{\phi}_{\pm}(0))$, where $a|_{\Sigma},\penalty 10000\ H|_{\Sigma}$ are related to $h_{ab},\penalty 10000\ \mathcal{K}_{ab}$, respectively. For a specific model of scalar fields or space-time dynamics, additional constraint equations may arise that must be satisfied on the Cauchy surface. Next, we discuss the requirement of imposing additional constraints on the scalar fields and their derivatives on $\Sigma$ to ensure that the scalar field closely mimics the behavior of cold dark matter (CDM).
Appendix C Additional Diagrams
C.1 Evolution of Cosmological Parameters when CDM decays into radiation